OpenStudy (yolo729):

Part A: If (2^6)^x = 1, what is the value of x? Explain your answer. Part B: If (5^0)6x = 1, what are the possible values of x? Explain your answer.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

anything to the power of zero is one

OpenStudy (mathmale):

Note that "x" in Part A is an exponent. Altho' it's not strictly necessary, you might note that 2^6 is "2 to the sixth power."

OpenStudy (yolo729):

@3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (yolo729):

question at top

OpenStudy (3mar):

I saw it

OpenStudy (3mar):

For part one: it is an equation that x is in the power not in the base, so it is an exponent equation.

OpenStudy (3mar):

To solve such ones, you would better make all bases the same, then compare the indexes.

OpenStudy (3mar):

In our case: \[(2^6)^x=1\] that means that: \[2^{6x}=1\] and you may know that any base (number) raised to the power zero equals 1 so we can use that fact to do the following \[1=2^0\]

OpenStudy (mathmale):

An alternative approach would be to apply the natural log operator to both sides of the first given equation. Don't know whether or not you've reached this point yet, but finding a "log" is the "inverse operation" of "exponentiation." Hope this is a review for you:\[\log 10^x = x\] \[10^{\log x}=x\]

OpenStudy (3mar):

So we can conclude that: \[2^{6x}=2^0\] the same bases >> equal indexes \[6x=0>>>x=0\] Got the first part?

OpenStudy (yolo729):

yes

OpenStudy (mathmale):

Your result?

OpenStudy (3mar):

Can you show me your work for the second?

OpenStudy (mathmale):

As 3mar has pointed out, (anything)^0 = 1.

OpenStudy (3mar):

Thank you for the medal! but let's hit the target firstly.

OpenStudy (yolo729):

ok

OpenStudy (3mar):

Can you share yours?

OpenStudy (yolo729):

it could be x=1

OpenStudy (yolo729):

or x=0

OpenStudy (yolo729):

this is for part b

OpenStudy (3mar):

But where your step to determine whichever 1 or 0?

OpenStudy (3mar):

How do you get 0 or 1?

OpenStudy (3mar):

@yolo729 Can you proceed?